Jillian is five years younger than Ciane. Four years ago, the sum of their ages was 25. How old are they now?

Found 3 solutions by josgarithmetic, amarjeeth123, MathTherapy:
Answer by josgarithmetic(39616)
Answer by amarjeeth123(569)

Let Jillian's present age be x.
Then Ciane's present age is (x+5).
Four years ago their ages were (x-4) and (x+1) respectively.
Four years ago the sum of their ages was (2x-3).
Plugging in the values we get,
2x-3=25
2x=28
x=14
Jillian's present age is 14 years.
Ciane's present age is 19 years.
Validation:
4 years back Jillian's age was 10 years.
4 years back Ciane's age was 15 years.
The sum of their ages 4 years back was 25 years.
Hence the solution is correct.

Answer by MathTherapy(10551)


Jillian is five years younger than Ciane. Four years ago, the sum of their ages was 25. How old are they now?
Let Ciane's age be C
Then Jillian's is: C - 5
Four years ago, Ciane and Jillian were C - 4, and C - 5 - 4, or C - 9, respectively
We then get: C - 4 + C - 9 = 25
2C - 13 = 25
2C = 38
Ciane's age, or 
Now, you should be able to find Jillian's age.